# A Classification-Driven Likelihood Ratio Method for Familial DNA Testing

### v1

@@ Abstract
Background Familial DNA search relies on statistical methods to infer genetic relationships, often using the likelihood ratio (LR) as the test statistic. This approach involves calculating the probability of observed DNA profiles under various scenarios, traditionally assuming a uniform genetic background for all individuals. However, this assumption is unrealistic in practice due to population substructure, where allele frequencies differ among subpopulations. To address this, modified LR-based approaches have been proposed, including methods such as calculating likelihood ratios using average allele frequencies (LRLAF) or the maximum, minimum, and average likelihood ratios (LRMAX, LRMIN, and LRAVG). Despite these developments, further advancements are needed to improve the testing of genetic relationships in the presence of subpopulation structures.
Results We propose a novel LR-based statistic designed to account for population substructure by incorporating a classification step to address nuisance parameters. Specifically, two DNA profiles with unknown subpopulation origins are simultaneously classified into one group. Following this classification, the likelihood ratio (LRCLASS) is applied. Our analysis demonstrates that the proposed LRCLASS statistic, when paired with Naive Bayes classification, exhibits higher statistical power than existing methods for testing full-sibling relationships in the Thai population. Additionally, we explored an alternative classification approach using multinomial logistic regression, which achieved performance comparable to the Naive Bayes method.
Conclusion The proposed LRCLASS statistic enhances the power of familial DNA searches by effectively addressing population substructure through a novel classification-based approach. This method provides a robust alternative to existing LR-based strategies and underscores the potential for integrating advanced classification techniques into genetic relationship testing. Future research could investigate other classification methods in conjunction with likelihood ratio calculations to further improve the accuracy and applicability of familial DNA analyses.
@@
familial DNA search, likelihood ratio, population substructure, classification
1 Background
@@ 1.1 Familial DNA Search
In forensic investigations, genetic material or DNA is one of the most common types of evidence collected from crime scenes. It can be extracted from various parts of the human body, such as hair, blood, sweat, or saliva. A typical question in such cases is whether a DNA sample from the crime scene matches any individual in a forensic database. If a match is found, the investigation becomes more straightforward, as the individual who left the sample can be identified. The results can then be submitted to the court for further proceedings. However, if the DNA sample from the crime scene does not perfectly match anyone in the database but shows some similarities, it may indicate a genetic relationship between two individuals, such as parent-child or full-sibling. This situation leads to familial DNA searches, which use statistical tools to infer familial relationships based on DNA evidence. Various strategies for familial searching have been proposed in the literature, ranging from basic methods that count the number of shared alleles to more advanced probabilistic approaches (e.g., Balding et al. 2013, Bieber et al. 2006, Ge et al. 2011, Ge & Budowle 2012, Kooakachai et al. 2019, Kruijver et al. 2014, Slooten & Meester 2014). Familial DNA searches have grown in prominence over the years. By 2014, at least 40 laboratories in the United States were conducting such searches, and approximately 75% of those not yet practicing this method were actively considering its implementation (Debus-Sherrill & Field, 2019). This approach is one of several indirect methods for identifying potential sources of forensic biological evidence, alongside techniques like Y-STR database searches and investigative genetic genealogy, as discussed by Ge and Budowle (2021). Despite its utility, familial DNA searching has sparked ethical concerns, particularly regarding the disproportionate representation of nonwhite individuals in databases and the potential for intrusive investigations targeting relatives of suspects. In response, U.S. states have adopted varied policies: some have formalized rules permitting or banning familial searches, while others allow follow-up investigations on accidental partial matches under regulatory oversight to mitigate ethical issues (Katsanis 2020).
@@
In statistical hypothesis testing for familial DNA searches, the likelihood ratio statistic is frequently employed to evaluate the probability of a pair of DNA profiles under two distinct relationship hypotheses. One scenario assumes the individuals are unrelated, while the other considers a specific familial relationship, such as parent-child or full-sibling. Over the years, the adoption of the likelihood ratio statistic for relatedness testing has grown significantly. A study on the New Zealand database, for instance, demonstrated that the likelihood ratio approach for parent-child and full-sibling tests outperformed the traditional method of counting matching alleles (Curran & Buckleton 2008). Building on this, a validation study further highlighted the effectiveness of a likelihood ratio-based method for identifying first-degree familial relationships using autosomal STR profiles in California’s 1,000,000-profile database (Myers et al. 2011). This approach was subsequently implemented in California’s Familial Search Program, where it played a pivotal role in solving the Grim Sleeper serial murder case. The likelihood ratio statistic is both straightforward and statistically robust, but it comes with notable limitations. A key drawback is its assumption that all individuals in the population share the same genetic background, meaning their genetic information is drawn from a single allele frequency distribution. However, this assumption is often unrealistic in forensic applications. Populations are more accurately modeled as having multiple genetic substructures, such as racial or ancestral groups, which can significantly influence allele frequencies. Failing to account for these differences can lead to inaccurate likelihood ratio calculations and erroneous conclusions, potentially resulting in miscarriages of justice (Balding & Nichols 1994). In this context, population substructure refers to the genetic variation that exists between distinct subgroups within a larger population. These variations must be incorporated into likelihood ratio calculations to ensure accuracy and fairness in forensic analyses.
@@
The critical role of subpopulation structure in forensic science is emphasized in numerous studies to ensure accurate DNA analysis and minimize errors. Factors such as geographical proximity, migration patterns, and cultural practices are shown to shape subpopulation structures, significantly influencing likelihood assessments in forensic investigations (Gallo et al. 1997). Allelic frequencies across diverse racial and ethnic databases have been examined, revealing substantial variations in probability calculations (Gallo et al. 1997). Differences in allele frequencies among ethnic groups, combined with genetic admixture, have been demonstrated to increase the risk of false positives or reduce the statistical power of genetic studies (Tian et al. 2008). Strategies to address population substructure in DNA database searches have been proposed, with simulations using ten-locus DNA profiles showing that even mild substructure can inflate evidence strength in 11.6% of cases (Spooner & Stockmarr 2019). A comprehensive overview of the importance of subpopulation structure has been provided, including discussions on foundational statistical concepts and recent advances, such as methods for ancestral classification using informative single nucleotide polymorphism markers (Jeong et al. 2021).
@@
When allele frequencies are available for each subpopulation, adjustments to the likelihood ratio formula have been proposed by various researchers. One notable approach, implemented by the Denver Crime Laboratory in Colorado, USA, in 2008, was inspired by the work of Prescutti et al. (2002). This method involves calculating likelihood ratios multiple times, applying a single population framework to each available allele frequency distribution. The highest value from these calculations, referred to as LRMAX, is then selected as the final likelihood ratio statistic. Alternatively, a different approach, introduced by Ge and Budowle (2012), suggests using the minimum likelihood ratio instead. This method argues that the minimum likelihood ratio provides a more accurate kinship index for the population associated with the forensic profile. Their research demonstrated that the minimum likelihood ratio consistently outperformed the maximum method in terms of accuracy. Two additional likelihood ratio-based statistics were introduced in 2019. The first, LRLAF, applies the classical likelihood ratio framework using average available allele frequencies, while the second, LRAVG, calculates the average of all likelihood ratios derived from each allele frequency distribution (Kooakachai et al. 2019). When these methods were compared for parent-child and full-sibling tests in the Colorado population, LRLAF was found to exhibit approximately 10% higher power for parent-child tests and 5% higher power for full-sibling tests compared to LRMAX. However, debate continues regarding the consistency of these methods, particularly when applied to populations outside of Colorado.
@@
Existing methods for familial DNA searches in the literature primarily rely on basic arithmetic, without incorporating advanced classification techniques. Instead of simply averaging likelihood ratios (LRAVG), averaging at the allelic level (LRLAF), or using the maximum (LRMAX) or minimum (LRMIN), machine learning methods can be employed to predict the subpopulation to which a DNA profile belongs, followed by a likelihood ratio calculation after classification. The use of classification has already been applied in some forensic subfields. For example, the application of penalized logistic regression in forensic toxicology was demonstrated for calculating likelihood ratios in a two-class classification setting (Biosa et al. 2020). The effectiveness of this framework was illustrated through a case study on alcohol biomarker data, highlighting its ability to distinguish chronic alcohol drinkers and its adaptability for use in forensic and analytical chemistry applications.
This study introduces a novel likelihood ratio-based statistic designed to assess the genetic relationship between two individuals. The proposed method leverages a Naive Bayes classification framework, wherein DNA samples are initially categorized into subpopulations, and the likelihood ratio is subsequently computed based on the resulting classifications. The efficacy of the proposed statistic will be rigorously evaluated through its Type I error rates and statistical power. Furthermore, an alternative classification approach employing multinomial logistic regression is introduced, enabling a comparative analysis of the classification performance between the Naive Bayes and multinomial logistic regression methodologies.
@@ 1.2 DNA Profile and Dataset Description
Humans differ due to the random combination of nucleobases—adenine (A), thymine (T), guanine (G), and cytosine (C)—inherited from both parents. The human genome, composed of about 3 billion nucleobases and approximately 20,000 genes across 23 pairs of chromosomes, contains all the genetic instructions for an individual. Genetic markers are variations in DNA sequences at specific positions on chromosomes. Among these markers, short tandem repeats (STRs) are especially useful in forensic science. STRs consist of DNA sequences, typically two to seven base pairs long, repeated multiple times. The number of repeats at each locus varies between individuals, allowing for identification. For example, a 28-base pair sequence like TCTATCTATCTATCTATCTATCTATCTA represents seven repeats of the four-base sequence TCTA, indicating an allele of length 7 at that position. Each person inherits one STR allele from each biological parent, and the number of repeats can differ between them. A genotype at a specific locus is represented by two alleles, such as 11,12, meaning the individual has 11 and 12 repeats of the TCTA sequence at that locus. The variation in STR repeat lengths among individuals makes STRs highly effective for forensic identification.
DNA databases, including government repositories, play a crucial role in criminal investigations by helping identify potential suspects. Since their inception, forensic databases have largely depended on STR markers due to their higher variability and stronger discriminatory ability compared to other genetic markers, which makes them vital for precise identification (Butler 2005).
@@
In this paper, a DNA profile is described as the collection of genotypes across multiple loci for an individual, with the number of loci used in forensic investigations differing by region. Here, Thailand is highlighted as a case study due to its recognized genetic substructure and documented genetic diversity, as supported by archaeological and linguistic research. Genetic differences have been observed between Thai-Malay Muslims, primarily in southern Thailand, and Thai-Buddhists from other regions (Kutanan et al. 2014a). Distinct genetic profiles have also been identified between Thai-northeastern populations and other groups, including those in northern Thailand, through mitochondrial DNA-HVR1 variation studies conducted in the same year (Kutanan et al. 2014b). More recently, significant genetic differences among subpopulations have been emphasized through a genome-wide analysis of 32 Thai subpopulations, underscoring the importance of careful participant selection for biomedical and clinical research (Kutanan et al. 2021). Although Thailand does not yet have a National DNA Database, a DNA database was established by the Central Institute of Forensic Science (CIFS) in 2004 to address crimes in the Deep South. This database, which has since expanded to include approximately 160,000 profiles of crime samples, individuals, and prisoners, utilizes autosomal and Y-STR markers. A 24-loci STR kit was introduced in 2016 to support both direct and familial matching (Boonderm et al. 2017).
@@
To simplify the analysis, four main Thai subpopulations, defined by geographical regions—North (NO), Northeastern (NE), Central (CT), and South (SO)—are considered in this study. Allele frequency distributions for these groups were derived from DNA samples of 929 unrelated Thai individuals analyzed across 15 loci (Shotivaranon et al. 2009). The sample sizes for the North, Northeastern, Central, and South subpopulations were 202, 304, 212, and 211, respectively. Based on the availability of data, 15 core loci are utilized in the study: D8S1179, D21S11, D7S820, CSF1P0, D3S1358, TH01, D13S317, D16S539, D2S1338, D19S433, vWA, TPOX, D18S51, D5S818, and FGA.
2 Methods
@@
Calculating the probabilities in the numerator and denominator of $\Lambda$ requires knowledge of the allele frequency distributions. Let $f_{X_{1}}$ and $f_{X_{2}}$ represent the allele frequency distributions for the populations to which $X_{1}$ and $X_{2}$ belong, respectively. These distributions provide the frequencies of each allele, allowing the probabilities to be fully determined. That is, $f_{X_{1}}$ and $f_{X_{2}}$ are essential inputs for calculating the likelihood ratio, serving as nuisance parameters. Additionally, the probabilities are calculated separately for each locus and then combined by multiplying them, based on the assumption that loci are independent. The assumption of independence between loci requires that the loci are not genetically linked and are in linkage equilibrium. However, some core loci do not fully satisfy this assumption. For instance, the loci vWA and D12S391 are located on the same chromosome, approximately 6.3 megabases (Mb) apart. Ideally, genetic markers on the same chromosome should be separated by more than 50 Mb to ensure independence between genotypes at the two loci (O’Connor et al. 2011). Studies have indicated that independence between vWA and D12S391 cannot always be assumed at the individual level. Nevertheless, at the population level, there is no significant evidence of correlation between genotypes at the core loci, including these two loci (Budowle et al. 2011, Ge et al. 2012). For simplicity in our calculations, we proceed under the assumption of independence between loci.

### v2

# Incorporating Naive Bayes Classification to Address Subpopulation Structure in Familial DNA Search

@@ Abstract
Familial DNA search evaluates the genetic relatedness of two individuals by comparing the likelihood of their observed DNA profiles under two competing hypotheses—the null hypothesis that the individuals are unrelated and the alternative hypothesis that they are related—most commonly through the likelihood ratio (LR). Standard LR-based approaches typically assume a uniform genetic background; however, this assumption is rarely valid due to population substructure, where allele frequencies vary among subpopulations and can bias relationship inference. Existing modifications—such as LR calculations based on average allele frequencies (LRLAF) and strategies using maximum, minimum, or average likelihood ratios (LRMAX, LRMIN, LRAVG)—help mitigate these challenges but remain limited in their ability to fully address subpopulation differences. This study introduces a new LR-based statistic, LRCLASS, which incorporates a classification step using the Naive Bayes classifier to account for nuisance parameters associated with unknown subpopulation origins. In LRCLASS, the two DNA profiles being compared are jointly assigned to a subpopulation group via Naive Bayes before LR computation. Empirical evaluations using Thai population data show that LRCLASS achieves higher statistical power for detecting full-sibling relationships than existing LR-based methods. We further assessed multinomial logistic regression as an alternative classifier and found its performance comparable to that of Naive Bayes, suggesting flexibility in classifier choice. Overall, integrating the Naive Bayes classifier with LR computation offers a robust strategy for addressing population substructure in familial DNA search and highlights the broader potential of combining supervised learning techniques with forensic statistical methodologies to enhance the accuracy and reliability of genetic relationship testing.
@@
familial DNA search, likelihood ratio, population substructure, naive Bayes classifier, classification, kinship inference
1 Introduction
@@ 1.1 Familial DNA Search
In forensic investigations, genetic material or DNA is one of the most common types of evidence collected from crime scenes. It can be extracted from various parts of the human body, such as hair, blood, sweat, or saliva. A typical question in such cases is whether a DNA sample from the crime scene matches any individual in a forensic database. If a match is found, the investigation becomes more straightforward, as the individual who left the sample can be identified. The results can then be submitted to the court for further proceedings. However, if the DNA sample from the crime scene does not perfectly match anyone in the database but shows some similarities, it may indicate a genetic relationship between two individuals, such as parent-child or full-sibling. This situation leads to familial DNA searches, which use statistical tools to infer familial relationships based on DNA evidence.
Various strategies for familial searching have been proposed in the literature, ranging from basic methods that count the number of shared alleles to more advanced probabilistic approaches (e.g., Balding et al. 2013, Bieber et al. 2006, Ge et al. 2011, Ge and Budowle 2012, Kooakachai et al. 2019, Kruijver et al. 2014, Slooten and Meester 2014). Familial DNA searches have grown in prominence over the years. By 2014, at least 40 laboratories in the United States were conducting such searches, and approximately 75% of those not yet practicing this method were actively considering its implementation (Debus-Sherrill and Field, 2019). Familial DNA search is one of several indirect methods used to identify potential sources of forensic biological evidence, along with approaches such as Y-STR database searches and investigative genetic genealogy, as noted by Ge and Budowle (2021). Although useful, this technique has raised ethical concerns, especially because nonwhite individuals are often overrepresented in forensic databases, increasing the likelihood that their families may be subjected to unnecessary or intrusive investigations. As a result, U.S. states have taken different regulatory approaches: some have created formal policies that either permit or prohibit familial DNA searches, while others allow investigations to proceed when partial matches arise accidentally, but only under specific oversight intended to reduce ethical risks (Katsanis, 2020).
@@
In hypothesis testing for familial DNA searches, the likelihood ratio is frequently used to evaluate the probability of a pair of DNA profiles under two distinct relationship hypotheses. One scenario assumes that the individuals are unrelated, while the other considers a specific familial relationship, such as parent–child or full-sibling. Over the years, the adoption of the likelihood ratio statistic for relatedness testing has grown significantly. For example, a study on the New Zealand database showed that the likelihood ratio approach for parent–child and full-sibling tests outperformed the traditional method of counting matching alleles (Curran and Buckleton, 2008). Building on this, a validation study further demonstrated the effectiveness of a likelihood ratio–based method for identifying first-degree familial relationships using autosomal STR profiles in California’s 1,000,000-profile database (Myers et al., 2011).
Although the likelihood ratio statistic is both straightforward and statistically robust, it has notable limitations. A key drawback is its assumption that all individuals in the population share the same genetic background, meaning their genetic information is drawn from a single allele frequency distribution. In forensic applications, however, this assumption is often unrealistic. Populations are more accurately modeled as having multiple genetic substructures, such as racial or ancestral groups, which can significantly influence allele frequencies. Failing to account for these differences can lead to inaccurate likelihood ratio calculations and erroneous conclusions, potentially resulting in miscarriages of justice, as suggested by Balding and Nichols (1994). In this context, population substructure refers to genetic variation that exists between distinct subgroups within a larger population. These variations must be incorporated into likelihood ratio calculations to ensure accuracy and fairness in forensic analyses.
@@
The critical role of subpopulation structure in forensic science has therefore been emphasized in numerous studies to ensure accurate DNA analysis and minimize errors. Factors such as geographical proximity, migration patterns, and cultural practices shape subpopulation structures, directly influencing likelihood assessments in forensic investigations (Gallo et al., 1997). Analyses of allelic frequencies across racial and ethnic databases reveal substantial variations in probability calculations, and differences in allele frequencies—combined with genetic admixture—have been shown to increase the risk of false positives or reduce the statistical power of genetic studies (Tian et al., 2008). Strategies to address population substructure in DNA database searches have also been proposed; for instance, simulations using ten-locus DNA profiles show that even mild substructure can inflate evidence strength in 11.6% of cases (Spooner and Stockmarr, 2019). A comprehensive overview of the importance of subpopulation structure has further been provided by Jeong et al. (2021), who discuss foundational statistical concepts and recent advances, including methods for ancestral classification using informative single nucleotide polymorphism markers.
@@
When allele frequencies are available for each subpopulation, several adjustments to the likelihood ratio formula have been proposed. One notable approach, implemented by the Denver Crime Laboratory in Colorado, USA, in 2008, was inspired by the work of Prescutti et al. (2002). This method involves calculating likelihood ratios multiple times, applying a single-population framework to each available allele frequency distribution. The highest value from these calculations, referred to as LRMAX, is then selected as the final likelihood ratio statistic. Alternatively, Ge and Budowle (2012) proposed using the minimum likelihood ratio instead, arguing that it provides a more accurate kinship index for the population associated with the forensic profile. Their research demonstrated that the minimum likelihood ratio consistently outperformed the maximum method in terms of accuracy. Two additional likelihood ratio–based statistics were introduced in 2019: LRLAF, which applies the classical likelihood ratio framework using average available allele frequencies, and LRAVG, which calculates the average of all likelihood ratios derived from each allele frequency distribution (Kooakachai et al., 2019). When these methods were compared for parent–child and full-sibling tests in the Colorado population, LRLAF showed approximately 10% higher power for parent–child tests and 5% higher power for full-sibling tests compared to LRMAX. However, debate continues regarding the consistency and generalizability of these methods, particularly when applied to populations outside of Colorado.
@@
Existing methods for familial DNA searches in the literature primarily rely on basic arithmetic and do not incorporate advanced classification techniques. Rather than simply averaging likelihood ratios (LRAVG), averaging at the allelic level (LRLAF), or using the maximum (LRMAX) or minimum (LRMIN), machine learning approaches can be used to predict the subpopulation to which a DNA profile belongs, followed by likelihood ratio calculation after classification. The use of classification has already been applied in certain forensic subfields. For example, penalized logistic regression was employed in forensic toxicology to calculate likelihood ratios in a two-class classification setting (Biosa et al., 2019). The effectiveness of this framework was demonstrated through a case study on alcohol biomarker data, highlighting its ability to distinguish chronic alcohol drinkers and its potential adaptability for broader forensic and analytical chemistry applications.
This study introduces a novel likelihood ratio–based statistic designed to evaluate the genetic relationship between two individuals while explicitly accounting for population substructure. The approach begins by classifying each DNA profile into a subpopulation using a Naive Bayes classifier, after which the likelihood ratio is computed based on the predicted subpopulations. This framework aims to improve the accuracy and robustness of familial inference compared with existing LR-based approaches. In addition, we implement an alternative classification method using multinomial logistic regression, enabling a systematic comparison of classification performance. Section 2 describes how DNA profiles are stored and presents both existing and proposed statistics. Section 3 outlines the algorithms used to evaluate statistical power and compare classification strategies. Section 4 presents the results, and Sections 5 and 6 provide discussion and conclusions.
2 Methods
@@ 2.1 DNA Profile and Dataset Description
Humans differ due to the random combination of nucleobases—adenine (A), thymine (T), guanine (G), and cytosine (C)—inherited from both parents. The human genome, composed of about 3 billion nucleobases and approximately 20,000 genes across 23 pairs of chromosomes, contains all the genetic instructions for an individual. Genetic markers are variations in DNA sequences at specific positions on chromosomes. Among these markers, short tandem repeats (STRs) are especially useful in forensic science. STRs consist of DNA sequences, typically two to seven base pairs long, repeated multiple times. The number of repeats at each locus varies between individuals, allowing for identification. For example, a 28-base pair sequence like TCTATCTATCTATCTATCTATCTATCTA represents 7 repeats of the four-base sequence TCTA, indicating an allele of length 7 at that position. Each person inherits one STR allele from each biological parent, and the number of repeats can differ between them. A genotype at a specific locus is represented by two alleles, such as 11,12, meaning the individual has 11 and 12 repeats of the TCTA sequence at that locus. The variation in STR repeat lengths among individuals makes STRs highly effective for forensic identification.
@@
In this paper, a DNA profile is defined as the collection of genotypes across multiple loci for an individual, with the number of loci used in forensic investigations varying by region. Thailand is highlighted as a case study due to its well-documented genetic substructure and diversity, supported by archaeological and linguistic research. Genetic differences have been observed between Thai-Malay Muslims, primarily in southern Thailand, and Thai-Buddhists from other regions (Kutanan et al., 2014a). Distinct genetic profiles have also been identified between northeastern Thai populations and other groups, including those in northern Thailand, through mitochondrial DNA HVR1 variation studies conducted in the same year (Kutanan et al., 2014b). More recently, genome-wide analyses of 32 Thai subpopulations have further emphasized significant genetic differences, highlighting the importance of careful participant selection in biomedical and clinical research (Kutanan et al., 2021).
Although Thailand does not yet have a National DNA Database, a DNA database was established by the Central Institute of Forensic Science (CIFS) in 2004 to address crimes in the Deep South. This database has since expanded to include approximately 160,000 profiles from crime samples, individuals, and prisoners, utilizing both autosomal and Y-STR markers. In 2016, a 24-loci STR kit was introduced to support both direct and familial matching (Boonderm et al., 2017). DNA databases, including government repositories, play a crucial role in criminal investigations by helping identify potential suspects. STR markers are particularly important because of their high variability and strong discriminatory power, which makes them essential for precise identification (Butler, 2005).
@@
To perform analyses relevant to Thailand’s genetic diversity, this study focuses on four major Thai subpopulations defined by geographical regions: North (NO), Northeastern (NE), Central (CT), and South (SO). Allele frequency distributions for these groups were obtained from DNA samples of 929 unrelated Thai individuals, analyzed across 15 loci (Shotivaranon et al., 2009). The sample sizes for the North, Northeastern, Central, and South subpopulations were 202, 304, 212, and 211, respectively. Based on the available data, the 15 core loci used in this study are: D8S1179, D21S11, D7S820, CSF1P0, D3S1358, TH01, D13S317, D16S539, D2S1338, D19S433, vWA, TPOX, D18S51, D5S818, and FGA.
2.2 Hypothesis Test
@@
Calculating the probabilities in the numerator and denominator of $\Lambda$ requires knowledge of the allele frequency distributions. Let $f_{X_{1}}$ and $f_{X_{2}}$ represent the allele frequency distributions for the populations to which $X_{1}$ and $X_{2}$ belong, respectively. These distributions provide the frequencies of each allele, allowing the probabilities to be fully determined. In other words, $f_{X_{1}}$ and $f_{X_{2}}$ are essential inputs for calculating the likelihood ratio, serving as nuisance parameters. Additionally, the probabilities are calculated separately for each locus and then combined by multiplying them, based on the assumption that loci are independent. The assumption of independence between loci requires that the loci are not genetically linked and are in linkage equilibrium. In fact, some core loci do not fully satisfy this assumption. For instance, the loci vWA and D12S391 are located on the same chromosome, approximately 6.3 megabases (Mb) apart. Ideally, genetic markers on the same chromosome should be separated by more than 50 Mb to ensure independence between genotypes at the two loci (O’Connor et al., 2011). Studies have indicated that independence between vWA and D12S391 cannot always be assumed at the individual level. Nevertheless, at the population level, there is no significant evidence of correlation between genotypes at the core loci, including these two loci (Budowle et al. 2011, Ge et al. 2012). For simplicity in our calculations, we proceed under the assumption of independence between loci.